OpenStudy (anonymous):

how to simplify sqrt(12^2+12^2x3) ?

OpenStudy (anonymous):

\[\sqrt{12^2+12^2\times 3} \]

OpenStudy (anonymous):

\[\sqrt{144 +432}= \sqrt{576} = \sqrt{24^2}=24\]

OpenStudy (anonymous):

what if u dunno 576=24^2

OpenStudy (anonymous):

the worked solns had sqrt(4x144)=24 how did they get that 4x144

OpenStudy (anonymous):

nvmm

OpenStudy (anonymous):

OK do you know what 576 is divisible by 4

OpenStudy (anonymous):

Yeah you could have done \[\sqrt{144*4}=2\sqrt{144} = 2*12= 24\]

OpenStudy (anonymous):

yeah my bad ..

OpenStudy (anonymous):

is there a way to find highest perfect square ? cant remmeber l..

OpenStudy (anonymous):

You need to remember these numbers 1, 4, 9, 16, 25, 36, 49, 64, 81, 100, 121, 144, 196, 225.

OpenStudy (anonymous):

hm cant remember but i think teacher showed something like keep halving until u find squares.. 3 years ago ..

OpenStudy (anonymous):

like 576 = 2x288= 4x144

OpenStudy (anonymous):

OH yeah that could work

OpenStudy (anonymous):

oh is it just breaking it down into its factors .. xd

OpenStudy (anonymous):

Thats it =D

OpenStudy (anonymous):

thankss